# Revolutionizing Finance with LLMs: An Overview of Applications and Insights

### v2

@@ 3.1.1 Quantitative Trading
Quantitative trading has traditionally relied on mathematical and statistical models to drive investment decisions, often centering around historical data and predefined algorithmic strategies. This approach, while effective in certain market conditions, faces challenges in dynamic and complex market environments. Traditional quantitative models can struggle to adapt quickly to new information, particularly when it comes to unstructured data sources like news articles, social media, and financial reports. These sources contain valuable sentiment and opinion-based information that standard quantitative methods may overlook [66, 135].
@@
The integration of LLMs into quantitative trading strategies represents a significant advancement in the field. It allows for a more holistic approach to investment decisions, one that combines the precision of quantitative models with the nuanced understanding of market sentiments. This synergy not only enhances the robustness of trading strategies but also provides a competitive edge in rapidly changing market conditions. As the financial markets continue to evolve, the role of LLMs in quantitative trading is poised to become increasingly vital, marking a paradigm shift in how investment decisions are made. [123]
@@ 3.1.2 Portfolio Optimization
Traditional portfolio optimization, grounded in the principles of modern portfolio theory, seeks to balance risk against return, typically relying on historical market data and statistical analysis [98]. This approach, while systematic, often encounters challenges in rapidly evolving markets where historical data may not adequately predict future trends. Additionally, traditional models may not fully account for complex, real-world factors like geopolitical events or sudden market shifts, potentially leading to suboptimal asset allocations [3].
The integration of LLMs in portfolio optimization heralds a significant advancement in addressing these challenges. LLMs excel in processing and analyzing vast amounts of unstructured data, including market reports, news articles, and financial statements, providing deeper insights and supplementary analysis crucial for risk assessment. These models can uncover subtle market sentiments and emerging trends hidden in textual data, offering a more nuanced view of potential risks and opportunities. By augmenting quantitative data with qualitative insights derived from LLMs, investors can achieve a more holistic approach to portfolio optimization. This synergy not only enhances the robustness of traditional models but also equips investors with a more adaptive and informed strategy in the face of market uncertainties. [45]
@@ 3.1.3 Robo-advisors
The essence of robo-advisors’ appeal lies in their computational power, which allows them to tailor portfolios to the individual user’s circumstances, taking into account market dynamics and personal risk preferences. The LLMs is critical in this context, parsing extensive data sets to discern complex financial market patterns, allowing robo-advisors to provide informed investment guidance. Throughout the investment cycle, they continuously monitor portfolio performance, adjusting the balance between expected returns and user-defined risk thresholds [16].
A key benefit of robo-advisors is their ability to flexibly update investment strategies to reflect changes in the market, a flexibility often not available with traditional investing avenues. The enhanced flexibility can foster greater trust between financial advisors and their clients [9].While historical analysis has primarily focused on the algorithmic sophistication and legitimacy of robo-advisors [83], discussion is burgeoning about the psychological factors that guide individuals to use these AI platforms.
In a revealing study of the German robo-advisory market (covering approximately 78 assets), they examined approximately 243,000 portfolio pairs along with customer demographic data. The findings indicate that despite the high level of AI-driven sophistication, the personalization aspects of robo-advisory advice are currently limited [107]. Key factors that influence modern portfolio choices—such as the amount and nature (beta) of human capital or shadow assets—remain largely unresolved. Recommendations tend to cater to current investor biases or regulators’ views on portfolio allocation, which inadvertently limits the economic potential of robo-advisors while bolstering consumer confidence and ensuring regulatory sanctions. The renaissance of robo-advisory advice is highlighted by its tendency to eschew complex, customized strategies in favor of more broadly applicable investment principles, for reasons including explainability to the average user and the need for privacy and data security [7].
All in all, the integration of LLMs with robo-advisory services marks a quantum leap in the field of consulting. These AI-centric platforms will revolutionize investing and wealth management by connecting intricate financial acumen with the understanding of ordinary investors, although there is caution about their current scope for customization [90] [96].
@@ 3.2.1 Merge and Acquisition Forecasting
In Mergers and Acquisitions (M&A) forecasting, NLP offers pivotal tools for mining and interpreting vast arrays of textual data [130, 120]. LLMs can adeptly analyze financial reports, news articles, and press releases to unearth underlying trends or strategic shifts that may hint at forthcoming M&A activities [129].
@@ 3.2.2 Insolvency Forecasting
For insolvency forecasting, language models can analyze a myriad of textual sources to gauge a company’s financial health accurately. By evaluating financial disclosures, news articles, and statements from corporate leaders, these models can detect early signs of financial distress [19].
To complement traditional numerical modeling, these models integrate textual analysis from various reports and news sources into bankruptcy prediction models, enhancing their predictive accuracy [2]. The sentiment and tone in corporate communications and financial discussions can be meticulously analyzed and reveal early warnings of a company’s deteriorating financial situation.
Moreover, a critical examination of regulatory filings through NLP can reveal subtle linguistic or disclosure patterns [51]. These patterns are frequently observed in the prelude to financial difficulties or impending insolvency, providing essential insights for stakeholders and investors.
@@ 3.2.3 Market Trend Forecast
Historically, the academic sphere has gravitated towards econometric models, such as ARIMA [94], and the finance industry has harnessed machine learning algorithms to predict stock movements. These methods, while effective to a degree, cannot often evolve rapidly with market conditions or explain their predictions transparently.
@@
NLP is increasingly being utilized in market forecasting to complement traditional quantitative analysis methods [81]. By processing and interpreting textual data from various sources such as news articles, financial reports, and social media, NLP provides valuable insights into market sentiment and trends [129]. This use of NLP helps in uncovering underlying patterns and correlations in market behavior that might not be immediately apparent from numerical data alone. In market forecasting, the ability of NLP to quickly analyze large volumes of text and extract relevant information plays a crucial role in making timely and informed predictions about market movements.
@@
  • Reducing Human Bias: By relying on data-driven insights [88], LLMs can help reduce human bias in market forecasting, leading to more objective and reliable predictions.
As the most powerful language model to date [113, 122], GPT-4 brings to the table its formidable prowess in processing vast datasets, extracting nuanced patterns, and synthesizing this information to generate predictions. Its capacity to parse through disparate data sources, including real-time financial news, historical price data, and burgeoning trends on social media platforms, allows it to construct a multi-faceted view of market conditions.
@@ 3.3.1 Credit Scoring
The significance of credit and risk assessments in the financial sector cannot be overstated, as these evaluations play a crucial role in maintaining financial stability. Credit assessment not only covers the possibility of assessing an individual borrower’s ability to repay, but also includes a variety of applications such as analyzing the risks of potential investments, evaluating the financial health of a company, and assisting financial institutions in making decisions about loan policies and interest rates. Traditionally, financial credit and risk assessment methods have predominantly been rule-based or reliant on machine learning algorithms [27]. However, these approaches exhibit limited flexibility across different tasks, often being tailor-made for specific objectives. Consequently, they struggle to generalize or integrate knowledge from diverse financial tasks. Moreover, such methods cannot leverage insights transferable across various financial activities. The advent of LLMs offers a promising avenue to transcend these limitations [132, 8, 141]. LLMs, with their prowess in multi-task learning and few-shot generalization, present an opportunity to redefine the landscape of financial assessments. Current research is exploring the potential of LLMs to identify correlations between disparate financial tasks and generalize across them. This capability marks a potential paradigm shift in credit and risk evaluation methodologies. The application of LLMs in this domain, however, is not without its challenges. For instance, the need to analyze tabular data, which contains symbolic information markedly different from the natural language data that LLMs are typically trained on, presents a significant hurdle. Additionally, ensuring that these models avoid biases in sensitive attributes such as age or gender is paramount. Despite these challenges, LLMs offer considerable advantages in processing and analyzing large volumes of textual data, such as loan applications and transaction histories. This capability enables them to extract valuable insights that can be instrumental in credit and risk analysis. By analyzing historical data and market trends, LLMs can assist analysts in gaining a deeper understanding of market dynamics and individual credit risks. Nevertheless, it’s important to recognize that the effectiveness of LLMs in credit scoring is still evolving. As the financial industry continues to integrate more advanced technological solutions, the role of LLMs in enhancing the accuracy and efficiency of credit and risk assessments will likely become more pronounced, heralding a new era in financial analytics.
@@ 3.3.2 ESG Scoring
Environmental, Social, and Governance (ESG) scoring is a critical metric in the contemporary business and investment landscape. It serves as a tool for evaluating a company’s commitment to environmental stewardship, social responsibility, and governance practices. ESG scores came about due to the financial world’s need to assess companies against these three criteria to identify the best performers in these aspects [23]. Private commercial firms whose primary clients are portfolio managers and other investors use tangible and intangible data to construct ESG scores to produce new data that meet investors’ needs [33].
A growing trend sees more companies being evaluated by sustainability rating agencies. The objective of these assessments is to generate relevant data for stakeholders interested in utilizing non-financial information about these companies. The information is particularly valuable for those looking to assess their investments or to develop investment portfolios based on sustainability criteria [37]. There are several common approaches for ESG scoring. Firstly, companies like Refinitiv and Bloomberg collect data from public sources but do not offer any value-adding input or scoring [146]. Secondly, ESG data providers combine both public and own-created data to evaluate ESG scores or ratings (e.g. MSCI). Thirdly, some companies focus on specialized ESG issues such as Carbon Disclosure Project.
@@ 3.3.3 Fraud Detection
As trade volume escalates and digital wallet technology advances, the realm of financial risk management is increasingly facing sophisticated high-tech criminal activities. A case in point: during the initial nine months of 2023, over 83,000 Americans fell prey to credit card fraud, leading to collective financial losses for the victims amounting to \$183 million  [^1] https://www.aura.com/learn/credit-card-scams. Consequently, the implementation of robust fraud detection applications is imperative to preserve the integrity of financial systems and safeguard both the institutions and their clients from financial losses [35]. Leveraging advanced reasoning and text mining capabilities, LLMs can significantly contribute to the identification of financial fraud in various domains including transactions, emails, profiles, contractors, and decentralized finance [84]. These LLMs serve as an initial filter, learning from customer transaction histories and detailed transaction information to isolate highly suspicious transactions from the billions processed [106], thereby substantially alleviating the manual labor burden involved in investigating vast quantities of transaction data. In this study, we use the PaySim simulates mobile money transactions dataset [82] to evaluate GPT-4’s effectiveness in detecting fraud.
@@ 3.3.4 Compliance Check
LLMs with zero-shot learning capabilities are becoming indispensable in the dynamic world of financial compliance, where regulations are in a constant state of flux. Zero-shot LLMs can adapt to new standards without the need for fine-tuning, which traditionally demands regular updates and a wealth of annotated data [142]. This characteristic is particularly beneficial for tasks such as audits, transaction monitoring, and reporting, as well as financial reporting and disclosure. In audits, zero-shot LLMs can immediately be deployed to parse and analyze documents, identifying inconsistencies and irregularities by understanding the underlying context, without the need for a model that is fine-tuned to specific audit criteria which may change over time [67]. This saves valuable time and resources in an environment where regulatory frameworks can shift unpredictably. For transaction monitoring and reporting, these LLMs excel at detecting anomalous patterns indicative of non-compliance or suspicious activities. They are capable of understanding transactional nuances and alerting to irregularities, all without prior fine-tuning to the specific rules that could be subject to change due to evolving regulations or market practices. When it comes to financial reporting and information disclosure, zero-shot LLMs ensure that disclosures align with current reporting standards such as IFRS, even as those standards are updated. They provide an agile response to changing requirements, highlighting discrepancies against the latest regulations without the need for retraining on new data sets. The zero-shot learning approach of LLMs not only bypasses the labor-intensive process of continuous model retraining but also mitigates the risk of outdated compliance checks in the rapidly changing financial landscape. This agility makes zero-shot LLMs a critical tool for financial institutions seeking to maintain compliance with the latest regulatory demands [23, 53].
@@ 3.4.1 Financial Education
GPT-4 is an advanced artificial intelligence language model developed by OpenAI that is capable of understanding and generating human-like natural language [26, 77, 86, 43, 63, 25, 68, 41, 15, 76, 111, 40, 59, 72]. This feature makes it a powerful tool for financial education, especially when it comes to explaining complex financial concepts, providing customized learning experiences, and enhancing user interaction [125].
@@
Secondly, GPT-4 has unique advantages in providing customized learning experience [78, 74, 115, 75, 70, 64, 29, 43]. It can adjust content and difficulty according to the user’s learning progress, interests and needs [40, 73, 44, 111, 143, 144, 62, 137]. For example, for beginners, GPT-4 can provide basic financial knowledge and concepts; for more experienced learners, it can provide more in-depth analysis and advanced topics. This personalized learning approach helps improve learning efficiency and user satisfaction.

### v3

@@ 3.1.1 Quantitative Trading
Quantitative trading has traditionally relied on mathematical and statistical models to drive investment decisions, often centering around historical data and predefined algorithmic strategies. This approach, while effective in certain market conditions, faces challenges in dynamic and complex market environments. Traditional quantitative models can struggle to adapt quickly to new information, particularly when it comes to unstructured data sources like news articles, social media, and financial reports. These sources contain valuable sentiment and opinion-based information that standard quantitative methods may overlook [liu2023fingpt, yu2023finme].
@@
The integration of LLMs into quantitative trading strategies represents a significant advancement in the field. It allows for a more holistic approach to investment decisions, one that combines the precision of quantitative models with the nuanced understanding of market sentiments. This synergy not only enhances the robustness of trading strategies but also provides a competitive edge in rapidly changing market conditions. As the financial markets continue to evolve, the role of LLMs in quantitative trading is poised to become increasingly vital, marking a paradigm shift in how investment decisions are made. [wang2023alphagpt]
@@ 3.1.2 Portfolio Optimization
Traditional portfolio optimization, grounded in the principles of modern portfolio theory, seeks to balance risk against return, typically relying on historical market data and statistical analysis [paulson1991makes]. This approach, while systematic, often encounters challenges in rapidly evolving markets where historical data may not adequately predict future trends. Additionally, traditional models may not fully account for complex, real-world factors like geopolitical events or sudden market shifts, potentially leading to suboptimal asset allocations [aligeopolitical].
The integration of LLMs in portfolio optimization heralds a significant advancement in addressing these challenges. LLMs excel in processing and analyzing vast amounts of unstructured data, including market reports, news articles, and financial statements, providing deeper insights and supplementary analysis crucial for risk assessment. These models can uncover subtle market sentiments and emerging trends hidden in textual data, offering a more nuanced view of potential risks and opportunities. By augmenting quantitative data with qualitative insights derived from LLMs, investors can achieve a more holistic approach to portfolio optimization. This synergy not only enhances the robustness of traditional models but also equips investors with a more adaptive and informed strategy in the face of market uncertainties. [jeong2024finetuning]
@@ 3.1.3 Robo-advisors
The essence of robo-advisors’ appeal lies in their computational power, which allows them to tailor portfolios to the individual user’s circumstances, taking into account market dynamics and personal risk preferences. The LLMs is critical in this context, parsing extensive data sets to discern complex financial market patterns, allowing robo-advisors to provide informed investment guidance. Throughout the investment cycle, they continuously monitor portfolio performance, adjusting the balance between expected returns and user-defined risk thresholds [caspigenerative].
A key benefit of robo-advisors is their ability to flexibly update investment strategies to reflect changes in the market, a flexibility often not available with traditional investing avenues. The enhanced flexibility can foster greater trust between financial advisors and their clients [baker2017regulating].While historical analysis has primarily focused on the algorithmic sophistication and legitimacy of robo-advisors [lourencco2020whose], discussion is burgeoning about the psychological factors that guide individuals to use these AI platforms.
In a revealing study of the German robo-advisory market (covering approximately 78 assets), they examined approximately 243,000 portfolio pairs along with customer demographic data. The findings indicate that despite the high level of AI-driven sophistication, the personalization aspects of robo-advisory advice are currently limited [scherer2023trust]. Key factors that influence modern portfolio choices—such as the amount and nature (beta) of human capital or shadow assets—remain largely unresolved. Recommendations tend to cater to current investor biases or regulators’ views on portfolio allocation, which inadvertently limits the economic potential of robo-advisors while bolstering consumer confidence and ensuring regulatory sanctions. The renaissance of robo-advisory advice is highlighted by its tendency to eschew complex, customized strategies in favor of more broadly applicable investment principles, for reasons including explainability to the average user and the need for privacy and data security [aw2023counteracting].
All in all, the integration of LLMs with robo-advisory services marks a quantum leap in the field of consulting. These AI-centric platforms will revolutionize investing and wealth management by connecting intricate financial acumen with the understanding of ordinary investors, although there is caution about their current scope for customization [maple2023ai] [nourallah2023one].
@@ 3.2.1 Merge and Acquisition Forecasting
In Mergers and Acquisitions (M&A) forecasting, NLP offers pivotal tools for mining and interpreting vast arrays of textual data [yang2021fact, visintin2023leveraging]. LLMs can adeptly analyze financial reports, news articles, and press releases to unearth underlying trends or strategic shifts that may hint at forthcoming M&A activities [yang2023fingpt].
@@ 3.2.2 Insolvency Forecasting
For insolvency forecasting, language models can analyze a myriad of textual sources to gauge a company’s financial health accurately. By evaluating financial disclosures, news articles, and statements from corporate leaders, these models can detect early signs of financial distress [chen2018models].
To complement traditional numerical modeling, these models integrate textual analysis from various reports and news sources into bankruptcy prediction models, enhancing their predictive accuracy [ahmadi2018towards]. The sentiment and tone in corporate communications and financial discussions can be meticulously analyzed and reveal early warnings of a company’s deteriorating financial situation.
Moreover, a critical examination of regulatory filings through NLP can reveal subtle linguistic or disclosure patterns [kim2021corporate]. These patterns are frequently observed in the prelude to financial difficulties or impending insolvency, providing essential insights for stakeholders and investors.
@@ 3.2.3 Market Trend Forecast
Historically, the academic sphere has gravitated towards econometric models, such as ARIMA [nelson1998time], and the finance industry has harnessed machine learning algorithms to predict stock movements. These methods, while effective to a degree, cannot often evolve rapidly with market conditions or explain their predictions transparently.
@@
NLP is increasingly being utilized in market forecasting to complement traditional quantitative analysis methods [lopez2023can]. By processing and interpreting textual data from various sources such as news articles, financial reports, and social media, NLP provides valuable insights into market sentiment and trends [yang2023fingpt]. This use of NLP helps in uncovering underlying patterns and correlations in market behavior that might not be immediately apparent from numerical data alone. In market forecasting, the ability of NLP to quickly analyze large volumes of text and extract relevant information plays a crucial role in making timely and informed predictions about market movements.
@@
  • Reducing Human Bias: By relying on data-driven insights [makridakis2023large], LLMs can help reduce human bias in market forecasting, leading to more objective and reliable predictions.
As the most powerful language model to date [sun2024trustllm, wang2024large], GPT-4 brings to the table its formidable prowess in processing vast datasets, extracting nuanced patterns, and synthesizing this information to generate predictions. Its capacity to parse through disparate data sources, including real-time financial news, historical price data, and burgeoning trends on social media platforms, allows it to construct a multi-faceted view of market conditions.
@@ 3.3.1 Credit Scoring
The significance of credit and risk assessments in the financial sector cannot be overstated, as these evaluations play a crucial role in maintaining financial stability. Credit assessment not only covers the possibility of assessing an individual borrower’s ability to repay, but also includes a variety of applications such as analyzing the risks of potential investments, evaluating the financial health of a company, and assisting financial institutions in making decisions about loan policies and interest rates. Traditionally, financial credit and risk assessment methods have predominantly been rule-based or reliant on machine learning algorithms [dastile2020statistical]. However, these approaches exhibit limited flexibility across different tasks, often being tailor-made for specific objectives. Consequently, they struggle to generalize or integrate knowledge from diverse financial tasks. Moreover, such methods cannot leverage insights transferable across various financial activities. The advent of LLMs offers a promising avenue to transcend these limitations [yoon2023design, babaei2023gpt, zheng2020effects]. LLMs, with their prowess in multi-task learning and few-shot generalization, present an opportunity to redefine the landscape of financial assessments. Current research is exploring the potential of LLMs to identify correlations between disparate financial tasks and generalize across them. This capability marks a potential paradigm shift in credit and risk evaluation methodologies. The application of LLMs in this domain, however, is not without its challenges. For instance, the need to analyze tabular data, which contains symbolic information markedly different from the natural language data that LLMs are typically trained on, presents a significant hurdle. Additionally, ensuring that these models avoid biases in sensitive attributes such as age or gender is paramount. Despite these challenges, LLMs offer considerable advantages in processing and analyzing large volumes of textual data, such as loan applications and transaction histories. This capability enables them to extract valuable insights that can be instrumental in credit and risk analysis. By analyzing historical data and market trends, LLMs can assist analysts in gaining a deeper understanding of market dynamics and individual credit risks. Nevertheless, it’s important to recognize that the effectiveness of LLMs in credit scoring is still evolving. As the financial industry continues to integrate more advanced technological solutions, the role of LLMs in enhancing the accuracy and efficiency of credit and risk assessments will likely become more pronounced, heralding a new era in financial analytics.
@@ 3.3.2 ESG Scoring
Environmental, Social, and Governance (ESG) scoring is a critical metric in the contemporary business and investment landscape. It serves as a tool for evaluating a company’s commitment to environmental stewardship, social responsibility, and governance practices. ESG scores came about due to the financial world’s need to assess companies against these three criteria to identify the best performers in these aspects [clement2022improving]. Private commercial firms whose primary clients are portfolio managers and other investors use tangible and intangible data to construct ESG scores to produce new data that meet investors’ needs [escrig2019rating].
A growing trend sees more companies being evaluated by sustainability rating agencies. The objective of these assessments is to generate relevant data for stakeholders interested in utilizing non-financial information about these companies. The information is particularly valuable for those looking to assess their investments or to develop investment portfolios based on sustainability criteria [friede2015esg]. There are several common approaches for ESG scoring. Firstly, companies like Refinitiv and Bloomberg collect data from public sources but do not offer any value-adding input or scoring [zumente2021esg]. Secondly, ESG data providers combine both public and own-created data to evaluate ESG scores or ratings (e.g. MSCI). Thirdly, some companies focus on specialized ESG issues such as Carbon Disclosure Project.
@@ 3.3.3 Fraud Detection
As trade volume escalates and digital wallet technology advances, the realm of financial risk management is increasingly facing sophisticated high-tech criminal activities. A case in point: during the initial nine months of 2023, over 83,000 Americans fell prey to credit card fraud, leading to collective financial losses for the victims amounting to \$183 million  [^1] https://www.aura.com/learn/credit-card-scams. Consequently, the implementation of robust fraud detection applications is imperative to preserve the integrity of financial systems and safeguard both the institutions and their clients from financial losses [feng2023empowering]. Leveraging advanced reasoning and text mining capabilities, LLMs can significantly contribute to the identification of financial fraud in various domains including transactions, emails, profiles, contractors, and decentralized finance [luo2023ai]. These LLMs serve as an initial filter, learning from customer transaction histories and detailed transaction information to isolate highly suspicious transactions from the billions processed [roy2018deep], thereby substantially alleviating the manual labor burden involved in investigating vast quantities of transaction data. In this study, we use the PaySim simulates mobile money transactions dataset [lopez2016paysim] to evaluate GPT-4’s effectiveness in detecting fraud.
@@ 3.3.4 Compliance Check
LLMs with zero-shot learning capabilities are becoming indispensable in the dynamic world of financial compliance, where regulations are in a constant state of flux. Zero-shot LLMs can adapt to new standards without the need for fine-tuning, which traditionally demands regular updates and a wealth of annotated data [zheng2023llm]. This characteristic is particularly beneficial for tasks such as audits, transaction monitoring, and reporting, as well as financial reporting and disclosure. In audits, zero-shot LLMs can immediately be deployed to parse and analyze documents, identifying inconsistencies and irregularities by understanding the underlying context, without the need for a model that is fine-tuned to specific audit criteria which may change over time [liu2023gpt]. This saves valuable time and resources in an environment where regulatory frameworks can shift unpredictably. For transaction monitoring and reporting, these LLMs excel at detecting anomalous patterns indicative of non-compliance or suspicious activities. They are capable of understanding transactional nuances and alerting to irregularities, all without prior fine-tuning to the specific rules that could be subject to change due to evolving regulations or market practices. When it comes to financial reporting and information disclosure, zero-shot LLMs ensure that disclosures align with current reporting standards such as IFRS, even as those standards are updated. They provide an agile response to changing requirements, highlighting discrepancies against the latest regulations without the need for retraining on new data sets. The zero-shot learning approach of LLMs not only bypasses the labor-intensive process of continuous model retraining but also mitigates the risk of outdated compliance checks in the rapidly changing financial landscape. This agility makes zero-shot LLMs a critical tool for financial institutions seeking to maintain compliance with the latest regulatory demands [clement2022improving, kojima2022large].
@@ 3.4.1 Financial Education
GPT-4 is an advanced artificial intelligence language model developed by OpenAI that is capable of understanding and generating human-like natural language [dai2023auggpt, liu2023deidgpt, ma2023impressiongpt, holmes2023evaluating, liao2023differentiate, dai2023adautogpt, liu2023summary, guan2023cohortgpt, cai2022coarse, liu2023pharmacygpt, shi2023mededit, gong2023evaluating, li2023artificial, liu2023context]. This feature makes it a powerful tool for financial education, especially when it comes to explaining complex financial concepts, providing customized learning experiences, and enhancing user interaction [wu2023bloomberggpt].
@@
Secondly, GPT-4 has unique advantages in providing customized learning experience [10.1007/978-3-031-45673-2_46, liu2023radiologyllama2, tang2023policygpt, liu2023radoncgpt, liuradiology, LIU2023100045, dou2023artificial, holmes2023evaluating]. It can adjust content and difficulty according to the user’s learning progress, interests and needs [gong2023evaluating, liu2023transformation, holmes2023benchmarking, shi2023mededit, zhong2023chatabl, zhou2023fine, liao2023mask, zhao2023ophtha]. For example, for beginners, GPT-4 can provide basic financial knowledge and concepts; for more experienced learners, it can provide more in-depth analysis and advanced topics. This personalized learning approach helps improve learning efficiency and user satisfaction.